# Non-negative Contrastive Learning

### v1

@@ 3.1 Benefits of Non-negativity: Consistency, Sparsity, and Orthogonality
Furthermore, NCL offers a robust solution to the rotation symmetry challenge associated with standard CL, as outlined in Theorem 1. Any rotation applied to features, especially sparse ones, often disrupts their non-negative nature. To exemplify, consider the feature matrix $I=\begin{pmatrix}1&0\\ 0&1\\ \end{pmatrix}.$ This matrix cannot undergo any non-permutative (i.e., non-permutative) rotation while preserving its non-negativity. As a consequence, NCL features are enforced to be aligned with the given axes. In particular, with non-negative constraints, activated samples along each feature dimension have high similarity and thus have similar semantics.
@@ 4.2 Optimal Representations
The following theorem states that when we have enough representation dimensions ($k=m$) [^2] For the ease of exposure, we consider $k=m$. If $k>m$, we can fill the rest of dimensions with all zeros, while if $k<m$, we can always tune $k$ to attain a lower loss. In practice, under standard CL training setting, there are some full zero dimensions (around 20%), so we have $k>m$. We omit zero dimensions for simplicity., the following solution $\phi(x)$ defined with $x$’s latent class information is an optimal solution of non-negative contrastive learning and thus enjoys nice explanations:
@@
where $[\pi_{1},\dots,\pi_{m}]$ is a random permutation of the subclass indices $[c_{1},\dots,c_{m}]$. [^3] Although we focus on the discrete latent class in the work, the discovered optimal solutions also apply to continuous latent variables, as long as the position generation still follows Assumption 1.
@@
Under the latent class assumption (Assumption 1) and choosing $k=m$, ϕ()italic-$\phi(\cdot)$ is a minimizer of the NCL objective Eq. (5), i.e., $\phi\in\operatorname*{arg\,min}{\mathcal{L}}_{\rm NCL}$.
Compared to CL with non-explanatory eigenvectors of $\bar{A}$ as the closed-form solution (HaoChen et al., 2021), NCL’s optimal features above are directly associated with the latent class probabilities, demonstrating clear interpretability. 1) Semantic Consistency. For each $c\in{\mathcal{C}}$, the activated samples with $\phi_{c}(x)>0$ indicates that ${\mathcal{P}}(c|x)>0$, meaning that $x$ indeed belongs to $c$ to some extent. Thus, activated samples along each dimension will have similar semantics. Instead, CL’s eigenvector features containing both positive and negative values do not have a clear semantic interpretation. 2) Sparsity. As each sample belongs to one or a few subclasses (Assumption 2), only one or a few elements in $\phi(x)$ are positive with ${\mathcal{P}}(c|x)>0$. Thus, $\phi(x)$ is sparse. In comparison, eigenvectors used by CL are usually dense vectors. 3) Orthogonality. Since feature dimensions in $\phi(x)$ indicate class assignments, low-class overlap (Assumption 2) implies that different feature dimensions will have very low correlation (nearly orthogonal):
@@ 5.1 Feature Selection
Here, we explore a simple strategy to rank non-negative features $f=[f_{1},\dots,f_{k}]$ according to their Expected Activation (EA) on the training data, $\operatorname{EA}(f_{i})=\mathbb{E}_{x}\tilde{f}_{i}(x)$, where $\tilde{f}(x)=f(x)/\|f(x)\|$ is normalized. Intuitively, higher EA indicates common high-level features that represent class semantics, and lower EA corresponds to rare and noisy features that are less useful. In our experiments, we select the top 256 features among all 512 features of projector outputs, and evaluate them with linear probing, image retrieval, and transfer learning on ImageNet-100 (details in Appendix G.4).
@@
All (512) 67.2 $\pm$ 0.2 68.0 $\pm$ 0.2 10.3 $\pm$ 0.1 11.4 $\pm$ 0.2 27.5 $\pm$ 0.1 30.6 $\pm$ 0.1
Random (256) 66.5 $\pm$ 0.3 65.3 $\pm$ 0.2 9.7 $\pm$ 0.1 7.8 $\pm$ 0.2 26.2 $\pm$ 0.4 24.7 $\pm$ 0.1
EA (256, w/o ReLU) 65.9 $\pm$ 0.2 66.2 $\pm$ 0.2 9.8 $\pm$ 0.2 8.4 $\pm$ 0.1 26.1 $\pm$ 0.1 25.3 $\pm$ 0.2
EA (256, w/ ReLU) 66.3 $\pm$ 0.2 67.5 $\pm$ 0.1 9.8 $\pm$ 0.1 11.2 $\pm$ 0.1 26.9 $\pm$ 0.2 30.5 $\pm$ 0.1
Results. Table 1 summarizes the results of feature selection experiments. First, the proposed EA criteria can obtain significant gains over random selection in NCL. With EA selection, NCL not only outperforms CL significantly, but also is almost “lossless”, since it attains comparable performance to using all features (difference $<0.5\%$). Thus, benefiting from the sparsity and disentanglement properties, NCL allows flexible feature selection that retains most of the original performance.
@@ 5.2 Feature Disentanglement
CL 0.35 $\pm$ 0.03 0.33 $\pm$ 0.01 0.25 $\pm$ 0.01 0.17 $\pm$ 0.01
NCL 3.31 $\pm$ 0.63 2.51 $\pm$ 0.37 0.69 $\pm$ 0.05 0.18 $\pm$ 0.01
Results. As shown in Table 2, NCL features show much better disentanglement than CL in all top-$k$ dimensions. The advantage is larger by considering the top features (e.g., 0.24 v.s. 0.7 SEPIN@$100$), since learned features also contain noisy dimensions. This verifies our identifiability result that with non-negativity constraints, NCL indeed brings better feature disentanglement on real-world data.
@@ 5.3 Downstream Generalization
Results. As shown in Table 3, NCL shows superior generalization performance than original contrastive learning across multiple real-world datasets. On average, it improves linear probing accuracy by $0.6\%$ and improves fine-tuning accuracy by $0.5\%$. Moreover, we observe from Table 3(b) that NCL shows significant advantages in the out-of-distribution generalization downstream tasks ($+1.8\%$ on average). Considering that NCL only brings minimal change to CL by adding a simple reparameterization with neglectable cost, the improvements are quite favorable. It also aligns well with the common belief that disentangled feature learning can yield representations that are more robust against domain shifts (Guillaume et al., 2019).
@@ 6 Extension to Broader Scenarios
Loss From Scratch Finetune CE 74.3 74.5 NCE 74.8 74.0 CE + MLP projector 76.3 74.1 NCE + MLP projector 77.3 76.9 Table 4: Test accuracy (%) of CE and NCE losses for supervised learning on ImageNet-100.
@@
Results. Figure 4 shows that non-negative training with NCE is much faster than conventional CE loss at the early stage (around 2x before the 40th epoch). After training converges, NCL also enjoys slightly better performance. As shown in Table 4, in terms of the final performance, NCE is helpful for training from scratch (0.5%). With the help of the MLP projector network, NCL attains even more significant gains by outperforming original CE by 3.0% and CE-with-projector by 1.0%. For finetuning, NCE alone is not helpful, but with MLP projector, NCE can attain a significant gain by improving 2.4% over CE, while the projector does not benefit the CE loss. These results show that although simple, non-negative training can be quite helpful for supervised tasks as well.
@@ Acknowledgement
Yisen Wang was supported by National Key R&D Program of China (2022ZD0160304), National Natural Science Foundation of China (62376010, 92370129), and Beijing Nova Program (20230484344).
@@ G.4 Experiment Details of Feature Selection
Transfer Learning on Selected Dimensions. In the downstream transfer learning tasks, we train a linear classifier on stylized ImageNet-100 (Geirhos et al., 2018) following the frozen representations with the default settings of linear probing. As shown in Table 1, the 256 dimensions with the largest EA values show superior performance, which implies that the estimated importance of NCL is robust and effective in downstream transfer learning tasks.
@@ G.5 Experiment Details of Downstream Classification
During the pretraining process, we utilize ResNet-18 (He et al., 2016) as the backbone and train the models on CIFAR-10, CIFAR-100 and ImageNet-100 (Deng et al., 2009). We pretrain the model for 200 epochs on CIFAR-10, CIFAR-100, ImageNet-100. We compare NCL and the original contrastive learning with the SimCLR (Chen et al., 2020) backbone. For CIFAR-10 and CIFAR-100, the projector is a two-layer MLP with hidden dimension 2048 and output dimension 256. And for ImageNet-100, the projector is a two-layer MLP with hidden dimension 4096 and output dimension 512. We pretrain the models with batch size 256 and weight decay 0.0001. When implementing NCL, we follow the default settings of SimCLR.

### v3

@@ 3.1 Benefits of Non-negativity: Consistency, Sparsity, and Orthogonality
Furthermore, NCL offers a robust solution to the rotation symmetry challenge associated with standard CL, as outlined in Theorem 1. Any rotation applied to features, especially sparse ones, often disrupts their non-negative nature. To exemplify, consider the feature matrix $I=\begin{pmatrix}1&0\\ 0&1\\ \end{pmatrix}.$ This matrix cannot undergo any non-permutative rotation while preserving its non-negativity. As a consequence, NCL features are enforced to be aligned with the given axes. In particular, with non-negative constraints, activated samples along each feature dimension have high similarity and thus have similar semantics.
@@ 4.2 Optimal Representations
The following theorem states that when the number of feature dimensions $k$ is as large as the number of latent classes $m$, the following solution $\phi(x)$ is an optimal solution to the NCL objective:
@@
where $[\pi_{1},\dots,\pi_{m}]$ is a random permutation of latent classes $[c_{1},\dots,c_{m}]$. [^2] For the ease of exposure, we consider $k=m$. If $k>m$, we can fill the rest of dimensions with all zeros, while if $k<m$, we can always tune $k$ to attain a lower loss. In practice, some feature dimensions are never activated, so we believe that generally we have $k>m$ and omit inactive dimensions for simplicity. Besides, although the discrete latent class is adopted here, the discovered optimal solutions also apply to continuous latent variables, as long as the position generation still follows Assumption 1. Nicely, $\phi(x)$ admits natural interpretability since it contains the posterior distribution along each dimension.
@@
Under the latent class assumption (Assumption 1) and choosing $k=m$, $\phi(\cdot)$ is a minimizer of the NCL objective Eq. (5), i.e., $\phi\in\operatorname*{arg\,min}{\mathcal{L}}_{\rm NCL}$.
Compared to CL with non-explanatory eigenvectors of $\bar{A}$ as the closed-form solution (HaoChen et al., 2021), NCL’s optimal features $\phi(x)$ are advantages in the following aspects. 1) Semantic Consistency. For each $c\in{\mathcal{C}}$, the activated samples with $\phi_{c}(x)>0$ indicates that ${\mathcal{P}}(c|x)>0$, meaning that $x$ indeed belongs to $c$ to some extent. Thus, activated samples along each dimension will have similar semantics. Instead, CL’s eigenvector features containing both positive and negative values do not have a clear semantic interpretation. 2) Sparsity. As each sample belongs to one or a few subclasses (Assumption 2), only one or a few elements in $\phi(x)$ are positive with ${\mathcal{P}}(c|x)>0$. Thus, $\phi(x)$ is sparse. In comparison, eigenvectors used by CL are usually dense vectors. 3) Orthogonality. Since feature dimensions in $\phi(x)$ indicate class assignments, that different feature dimensions have very low correlation when the latent class overlap $\varepsilon$ is small (Assumption 2):
@@ 5.1 Feature Selection
Here, we explore a simple strategy to rank non-negative features $f=[f_{1},\dots,f_{k}]$ according to their Expected Activation (EA) on the training data, $\operatorname{EA}(f_{i})=\mathbb{E}_{x}\tilde{f}_{i}(x)$, where $\tilde{f}(x)=f(x)/\|f(x)\|$ is normalized. Intuitively, higher EA indicates common high-level features that represent class semantics, and lower EA corresponds to rare and noisy features that are less useful. In our experiments, we select the top 512 features among all 2048 features of projector outputs, and evaluate them with linear probing, image retrieval, and transfer learning on ImageNet-100 (details in Appendix G.4).
@@
All (2048) 66.8 $\pm$ 0.2 68.9 $\pm$ 0.1 10.9 $\pm$ 0.2 14.2 $\pm$ 0.2 17.2 $\pm$ 0.1 19.9 $\pm$ 0.1
Random (512) 66.2 $\pm$ 0.1 (-0.6) 64.3 $\pm$ 0.2 (-5.6) 10 $\pm$ 0.1 (-0.9) 8.2 $\pm$ 0.1 (-6.0) 16.6 $\pm$ 0.2 (-0.6) 16.7 $\pm$ 0.1 (-3.2)
EA (512, w/o ReLU) 66.3 $\pm$ 0.2 (-0.5) 66.7 $\pm$ 0.1 (-2.2) 9.9 $\pm$ 0.21 (-0.9) 11.1 $\pm$ 0.2 (-3.1) 16.5 $\pm$ 0.3 (-0.7) 17.7 $\pm$ 0.2 (-2.2)
EA (512, w/ ReLU) (ours) 66.5 $\pm$ 0.1 (-0.3) 68.9 $\pm$ 0.3 (-0.0) 10.2 $\pm$ 0.2 (-0.7) 14.2 $\pm$ 0.2 (-0.0) 16.6 $\pm$ 0.3 (-0.6) 19.8 $\pm$ 0.1 (-0.1)
Results. Table 1 summarizes the results of feature selection experiments. First, when using all features, NCL achieves significant gains over CL on all three tasks. Meanwhile, when selecting 1/4 dimensions (512/2048), with EA criterion (w/ ReLU), NCL is almost “lossless”, since it attains comparable performance to using all features on all tasks. In comparison, CL features always suffer from a larger drop when using fewer features. Thus, benefiting from the sparsity and disentanglement properties, NCL allows flexible feature selection that retains most of the original performance.
@@ 5.2 Feature Disentanglement
CL 0.88 $\pm$ 0.08 0.79 $\pm$ 0.02 0.69 $\pm$ 0.01 0.47 $\pm$ 0.01
NCL 7.43 $\pm$ 0.15 5.93 $\pm$ 0.12 3.87 $\pm$ 0.04 0.48 $\pm$ 0.01
Results. As shown in Table 2, NCL features show much better disentanglement than CL in all top-$k$ dimensions. The advantage is larger by considering the top features (e.g., 0.69 v.s. 3.87 SEPIN@$100$), since learned features also contain noisy dimensions. This verifies our identifiability result that with non-negativity constraints, NCL indeed brings better feature disentanglement on real-world data.
@@ 5.3 Downstream Generalization
Results. As shown in Table 3, NCL shows superior generalization performance than original contrastive learning across multiple real-world datasets. On average, it improves linear probing accuracy by $0.6\%$ and improves fine-tuning accuracy by $0.9\%$. Moreover, we observe from Table 3(b) that NCL shows significant advantages in the out-of-distribution generalization downstream tasks ($+1.4\%$ on average). Considering that NCL only brings minimal change to CL by adding a simple reparameterization with neglectable cost, the improvements are quite favorable. It also aligns well with the common belief that disentangled feature learning can yield representations that are more robust against domain shifts (Guillaume et al., 2019).
@@ 6 Extension to Broader Scenarios
Loss From Scratch Finetune
CE 76.1 78.6
NCE 78.6 80.2
CE + MLP projector 78.4 81.1
NCE + MLP projector 79.2 82.0
Table 4: Test accuracy (%) of CE and NCE losses for supervised learning on ImageNet-100.
@@
Results. Figure 4 shows that non-negative training with NCE is much faster than conventional CE loss at the early stage (around 2x before the 40th epoch). After training converges, NCL also enjoys slightly better performance. As shown in Table 4, in terms of the final performance, NCE is helpful for training from scratch (2.5%). With the help of the MLP projector network, NCL also attains significant gains by outperforming original CE by 3.1% and CE-with-projector by 0.8%. For finetuning, with MLP projector, NCE can attain a significant gain by improving 3.4% over CE and 0.9 over CE-with projector. These results show that although simple, non-negative training can be quite helpful for supervised tasks as well.
@@ Acknowledgement
Yisen Wang was supported by National Key R&D Program of China (2022ZD0160300), National Natural Science Foundation of China (62376010, 92370129), and Beijing Nova Program (20230484344).
@@ G.4 Experiment Details of Feature Selection
Transfer Learning on Selected Dimensions. In the downstream transfer learning tasks, we train a linear classifier on stylized ImageNet-100 (Geirhos et al., 2018) following the frozen representations with the default settings of linear probing. As shown in Table 1, the 512 dimensions with the largest EA values show superior performance, which implies that the estimated importance of NCL is robust and effective in downstream transfer learning tasks.
@@ G.5 Experiment Details of Downstream Classification
During the pretraining process, we utilize ResNet-18 (He et al., 2016) as the backbone and train the models on CIFAR-10, CIFAR-100 and ImageNet-100 (Deng et al., 2009). We pretrain the model for 200 epochs on CIFAR-10, CIFAR-100, and 100 epochs for ImageNet-100. We compare NCL and the original contrastive learning with the SimCLR (Chen et al., 2020) backbone. For CIFAR-10 and CIFAR-100, the projector is a two-layer MLP with hidden dimension 2048 and output dimension 256. And for ImageNet-100, the projector is a two-layer MLP with hidden dimension 16384 and output dimension 2048. We pretrain the models with batch size 256 and weight decay 0.0001. When implementing NCL, we follow the default settings of SimCLR.